Suppose the XYZ Co. wants to sell preferred stock at $70 per share. A similar issue of preferred stock already outstanding has a price of $50 per share and offers a dividend of $1 every quarter. What dividend will XYZ Co. have to offer if the preferred stock is going to sell?

  1. $1.0 per share

  2. $1.2 per share

  3. $1.4 per share

  4. $1.8 per share

  5. $2.0 per share
Answer #1

Dividend Yield of stock trading at $50 = Dividend/ Price = 1/50= 0.02

Now using that dividend Yield , we will calculate the dividend of stock selling at $70= Price*dividend Yield

=70*0.02= $1.4 per share